Use the following data for Apple and Dell (in $ million):

Apple Dell
Profit (operating income) $6,275 $3,190
Sales revenue $32,479 $61,101
Investment (total assets) $32,460 $27,031


Required:

a) compute the ROI for each company.
Apple ROI=  % (if you get say 12.5%, enter 12.5, not 12.5% or 0.125)
Dell ROI=  %

b) use the DuPont method to decompose ROI into ROI = profit margin * asset turnover.
In other words, compute profit margin and asset turnover for each company:
Apple profit margin=  , asset turnover=  (Enter both as a fraction of 1, not as percentage. I.e., if the profit margin is 0.045 (4.5%) and asset turnover is 3.42 (342%), enter 0.045 and 3.42)
Dell profit margin=  , asset turnover=
If you multiply profit margin * asset turnover, you should get the ROI from part (a)

c) Which company has higher profit margin?(1=Apple, 2=Dell)

d) Which company has higher asset turnover? 1=Apple, 2=Dell)
Answer #1

a) ROI = Net income/Investment

Apple ROI = 6275/32460 = 19.3

Dell ROI = 3190/27031 = 11.8

b) Calculate following

Profit margin Turnover
Apple 6275/32479 = 0.193 32479/32460 = 1.00
Dell 3190/61101 = 0.052 61101/27031 = 2.26

c) Highest profit margin = Apple

d) Highest turnover = Dell